OpenStudy (anonymous):

can you please work this out convert the following decimal to percents 7.8 3/4

OpenStudy (paki):

is it 7.8(3/4)...?

OpenStudy (paki):

@applehead13

OpenStudy (anonymous):

yes

OpenStudy (paki):

ok... then it will be like this... 7.8*3/4 = 34.2/4 = (34.2/4) / 100 = 0.855%